If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Answer by stanbon(48558)

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
----------
Sally DATA:
Time = 4 hrs/job ; Rate = 1/4 job/hr.
-----------
John DATA:
Time = 6 hrs/job ; Rate = 1/6 job/hr.
-----------
Together DATA:
Time= x hrs/job ; Rate = 1/x job/hr.
----------
EQUATION:
rate + rate = rate together
1/4 + 1/6 = 1/x
(10/24)=1/x
Invert both sides to get:
x=24/10= 2.4 hours for them to do the job together.
Cheers,
Stan H.